OpenStudy (anonymous):

what does ln(2+x) +lnA simplify into? A is a constant

OpenStudy (anonymous):

when logarithms of the same base add, the terms inside multiply. \[\ln(a)+\ln(b)=\ln(a\times b)\]

OpenStudy (anonymous):

use that property to simplify

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw